Two polarizers are oriented at 58.0° to one another. Light polarized at a 29.0° angle to each

          polarizer passes through both. What reduction in intensity takes place?

          If instead the orientation of the polarizers is 40°to one another and that only 15% of the light gets through both of them, what was the initial polarization direction of the incident light?

          



1
Expert's answer
2021-03-02T07:09:05-0500

Initialy The light polarizez at 29^{\circ} from first poloroid. Let I_o be the unpolorized intensity of light.


Intensity of light after passing first poloroid

"I_1=I_ocos^{2}29^{\circ}\\I_1=0.765I_o"


Now the intensity of light after passing through second polarizers

"I_2=I_1cos^2(58^{\circ}-29^{\circ})\\I_2=I_1cos^229^{\circ}\\I_2=0.765I_o\times 0.765=0.585I_o"

Reduction in intensity of light= "\dfrac{I_o-0.585I_o}{I_o}=\dfrac{0.415I_o}{I_o}\times 100=41.5" %


Hence The intensity of light is reduced by 41.5%


If the orientation of the polarizers is 40°to one another and that only 15% of the light gets through both of them,


let the initial angle made by incident light be "\theta"


Then, After passing through second polarizer

"\Rightarrow\dfrac{15I_o}{100}=I_1cos^2(40-\theta)~~~~~~~-(1)"



Intensity of light passing from first polarizer

"I_1=I_ocos^2\theta~~~~~~~~-(2)"



From eqs.(1) and (2)-


"\dfrac{15}{100}=cos^2\theta cos^2(40-\theta)"


"\dfrac{1}{2}\times 2cos\theta cos(40^{\circ}-\theta)=\sqrt{0.15}"


"cos(\theta+40-\theta)+cos(\theta-40+\theta)=2\times 0.3872"


"\Rightarrow cos40+cos(2\theta-40)=0.7644\\\Rightarrow cos(2\theta-40)=0.7644-0.7664\\\Rightarrow 2\theta-40=cos^{-1}0\\\Rightarrow 2\theta=40+90\\\Rightarrow\theta=\dfrac{130}{2}"


"\Rightarrow \theta=65^{\circ}"


Hence the initial incidence angle is "65^{\circ}."